OpenStudy (anonymous):

I was trying to factor: f(t) = -16t^2 - 32t + 128 I got (-16 - 16)(t - 1)(t + 8). What did I do wrong?

OpenStudy (anonymous):

The first thing I did was split the middle term in half, so that I could factor by grouping.

OpenStudy (anonymous):

@mathmale @mathstudent55 @boldjon @ganeshie8 @sleepyjess @triciaal

OpenStudy (anonymous):

f(t) = -16t^2 - 32t + 128 = f(t) = -16t^2 - 16t - 16 + 128 Then i put them into groups f(t) = (-16t^2 - 16t) - (16 + 128)

OpenStudy (anonymous):

wait hold on for a minte

OpenStudy (marihelenh):

You could factor out the 16. That would simplify it enough.

OpenStudy (anonymous):

the next step that I did was change the + sign to a minus

OpenStudy (anonymous):

f(t) = (-16t^2 - 16t) + (-16 - 128)

OpenStudy (anonymous):

The GCF in the first equation is -16t^2 and in the second binomial it is -16

OpenStudy (anonymous):

So i divided the gcf out

OpenStudy (anonymous):

f(t) = -16t(t - 1) + -16(t - 128)

OpenStudy (anonymous):

Then I combined the terms (-16t - 16)(t - 1)(t - 8)

OpenStudy (marihelenh):

First, you can't split the middle term. It only has one variable, which means it can't be split. To split, each would have to take a variable or have no variable.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

hold on, then should i split the second term?

OpenStudy (marihelenh):

Going back to what I said, just factor out a -16.

OpenStudy (marihelenh):

Don't split any terms. This other way is super simple. You are over complicating this.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

f(t) = -16t2 - 32t + 128 / -16 f(t) = -16(t2 - 2t - 8)

OpenStudy (anonymous):

that all?

OpenStudy (marihelenh):

It would be +2t

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

but then it's done?

OpenStudy (marihelenh):

Then just simplify what is in parenthesis.

OpenStudy (anonymous):

ermm... how?

OpenStudy (marihelenh):

Do you know what FOIL is?

OpenStudy (anonymous):

FIRST INNER OUTER LAST VIDEO GAMES!

OpenStudy (anonymous):

yes, i do

OpenStudy (anonymous):

@marihelenh Can we hurry please? I'm in a serious rush lol

OpenStudy (marihelenh):

Yeah, it is pretty much just the reverse. You just have to solve for the equation.

OpenStudy (marihelenh):

A rush?

OpenStudy (anonymous):

how do I wait wut? i dont ge it

OpenStudy (marihelenh):

What are you in a rush for?

OpenStudy (anonymous):

unspecified, please continue :)

OpenStudy (marihelenh):

Well, I actually have to go to. Do you want me to explain later or just the answer?

OpenStudy (anonymous):

NO HELP HELP HELP HELP HELP PLEASE!

OpenStudy (anonymous):

DONT GO, STAY FOR 5 MORE MINUTES PLEASE!

OpenStudy (mathstudent55):

The first rule of factoring is to try to factor a common factor. \(f(t) = -16t^2 - 32t + 128 \) We see the coefficients are -16, 32, and 128. They are all divisible by 16. Let's factor out -16 from all terms: \(f(t) = -16(t^2 +2t -8) \)

OpenStudy (marihelenh):

He's got this! Sorry, but dinner's waiting!

OpenStudy (mathstudent55):

Now we need to factor the trinomial \(t^2 + 2t - 8\) It is a quadratic trinomial with a leading coefficient (the coefficient of the \(t^2\) term) of 1.

OpenStudy (anonymous):

@marihelenh alright, that's fine

OpenStudy (anonymous):

@mathstudent55 Yeah, I get it so far...

OpenStudy (mathstudent55):

Notice the numbers in blue and red. \(f(t) = -16(t^2 +\color{blue}{2}t \color{red}{-8})\) We split the trinomial into two binomials: \(f(t) = -16(t~~~~~~)(t~~~~~~)\) Now to complete the binomials, we need two numbers that multiply to \(\color{red}{-8}\) and add to \(\color{blue}{2}\).

OpenStudy (anonymous):

-2 and 4?

OpenStudy (mathstudent55):

Think of all pairs of numbers that multiply to -8. List them and list their sum. When you find a pair of such numbers that adds to 2, that is the pair of numbers you need to fill in the binomials above, and your factoring will be completed.

OpenStudy (anonymous):

-9 and 1, -4 - 4, -2 - 6, etc.

OpenStudy (anonymous):

-2 and 4 meet the demands, correct?

OpenStudy (boldjon):

yes correct

OpenStudy (anonymous):

hooray!

OpenStudy (anonymous):

@boldjon yea

OpenStudy (anonymous):

@mathstudent55 What's next?

OpenStudy (mathstudent55):

You already found that -2 * 4 = -8, and -2 + 4 = 2, so you have your numbers. Just as a note, you could have done this: List all pairs of numbers that multiply to -8. Follow each pair with the sum. -8 * 1 = -8; -8 + 1 = -7 -1 * 8 = -8; -1 + 8 = 7 -2 * 4 = -8; -2 + 4 = 2 <------- this is the only pair with the product and sum we need -4 * 2 = -8; -4 + 2 = -2

OpenStudy (mathstudent55):

Now that we have our two numbers, we just fill in the blanks I left above: \(f(t) = -16(t - 2)(t + 4) \)

OpenStudy (mathstudent55):

Now we look at each factor, and we see it is no longer factorable, so the original trinomial is factored.

OpenStudy (anonymous):

wait, where did you get the add to 2 and multiply to 8 anyways?

OpenStudy (anonymous):

nvm found out

OpenStudy (mathstudent55):

Look at the numbers in red and blue above.

OpenStudy (anonymous):

yea i found otut

OpenStudy (anonymous):

So the final answer is f(x) = 16(t−2)(t+4)?

OpenStudy (anonymous):

*-16

OpenStudy (mathstudent55):

Yes, with the -16.

OpenStudy (anonymous):

Wait, hold on. Basically I'm supposed to be finding the zeroe's and so I had to factor to do that. I'm supposed to find the x-intercept, where do I go from here?

OpenStudy (mathstudent55):

General Rule To factor a quadratic trinomial of the type (notice the coefficient of 1 for the \(x^2\) term): \(x^2+ax+b\) Find two numbers (let's call them \(p\) and \(q\)) whose product is \(b\) and whose sum is \(a\). Then the factoring is \((x+p)(x+q)\)

OpenStudy (anonymous):

??

OpenStudy (mathstudent55):

Ok. To find zeros, you need to set the function equal to zero. \(f(t) = -16t^2 - 32t + 128 = 0\) \(-16t^2 - 32t + 128 = 0\) \(-16(t - 2)(t + 4) = 0\) So far we have this with the factored polynomial.

OpenStudy (anonymous):

right

OpenStudy (mathstudent55):

Now we solve for t. \(-16(t - 2)(t + 4) = 0\)

OpenStudy (mathstudent55):

We have a product of factors equaling zero. A product can only equal zero if at least one of the factors equals zero. This is the zero product rule.

OpenStudy (mathstudent55):

\(-16(t - 2)(t + 4) = 0\) Since -16 is obviously not equal to zero, then the only factors that may equal zero are t - 2 and t + 4.

OpenStudy (anonymous):

alright...

OpenStudy (anonymous):

therefore, t = 2 and t = -4

OpenStudy (mathstudent55):

We set the two equations: \(t - 2 = 0\) or \(t + 4 = 0\) Can you solve those equations?

OpenStudy (anonymous):

correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?correct? correct?

OpenStudy (mathstudent55):

Oh, you did already.

OpenStudy (mathstudent55):

The zeros of the polynomial are t = 2 and t = -4. You are correct.

OpenStudy (anonymous):

but what happens to the -16?

OpenStudy (mathstudent55):

The zeros of a polynomial are the points where the polynomial crosses the x-axis.

OpenStudy (anonymous):

so where does teh -16 go?

OpenStudy (anonymous):

IBANGUOAENGPOAUIEGNEAOSIGNAOEGNOPAEGOAWIEGNAEOMF ^Tantrum

OpenStudy (mathstudent55):

The -16: We had \(-16(t - 2)(t +4)= 0\) Divide both sides by -16: \(\dfrac{\cancel{-16}(t - 2)(t + 4)}{\cancel{-16}~~1} = \dfrac{0}{-16}\) The -16's cancel out on the left side. On the right side, 0/(-16) is simply 0. We end up with only \((t - 2)(t + 4) = 0\) Which has roots (or zeros) 2 and -4. The -16 does not gives us any more zeros.

OpenStudy (anonymous):

I don't really get it... but okaay

OpenStudy (anonymous):

anyways, thanks for helping!

OpenStudy (mathstudent55):

The -16 was simply canceled out by dividing both sides by -16.

OpenStudy (anonymous):

Can you give a medal to @marihelenh , she/he helped too.

OpenStudy (mathstudent55):

Sure.

OpenStudy (mathstudent55):

The zeros are the x-intercepts.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

@mathstudent55 Can I tag you in another question?